OpenStudy (anonymous):

Solve the variable in this equation. 5x/7 - 1/14 = 9/28

OpenStudy (anonymous):

Someone help me solve this ASAP!

OpenStudy (campbell_st):

start by adding 1/14 to both sides of the equation... then multiply every term by 7 lastly divide by both sides by 5

OpenStudy (anonymous):

After dividing both sides by 5 what would the answer be?

OpenStudy (nurali):

\[\frac{ 5x }{ 7 }-\frac{ 1 }{ 14 }=\frac{ 9 }{ 28 }\] \[\frac{ 5x }{ 7 }=\frac{ 9 }{ 28 }+\frac{ 1 }{ 14 }\]

OpenStudy (campbell_st):

well tell me what you have...

OpenStudy (campbell_st):

are you doing a quiz...?

OpenStudy (nurali):

\[\frac{ 5x }{ 7 }=\frac{ 11 }{ 28 }\]

OpenStudy (nurali):

Multiply both side by 7 and then divide by 5

OpenStudy (anonymous):

My answers are coming out way different than the answer options it's giving me.

OpenStudy (campbell_st):

well what answers have you been getting...?